What is the pH of a buffer that is 0.300 M HF (Kg 6.8x104) and 0.500 M KF? 9.
Answer #1

Applying Handerson Hasselbach equation

pH = pKa + log [salt] / [acid]

Ka = 6.8 * 10^-4

hence pKa = -log Ka = -log (6.8 * 10^-4)=3.17

pH = 3.17 + log [0.5] / [[0.3]

pH = 3.39

PLEASE LEAVE THE POSITIVE FEEDBACK IF YOU LIKE THE SOLUTION